I need help solving this the vertex of the graph of this function:
f(x)=-4x^2+24x-40
I also need the axis of the graph of function for:
f(x)=4x^2-32x+27
As well as the quadratic function in vertex form of:
y=x^2+8x
Thanks for any help

Answer by lwsshak3(6505)

Standard form of equation for a parabola: y=A(x-h)^2+k, (h,k) being the (x,y) coordinates of the vertex
..
I need help solving this the vertex of the graph of this function:
f(x)=-4x^2+24x-40
complete the square:
y=-4(x^2-6x+9)-40+36
y=-4(x-3)^2-4
vertex: (3,-4)
..
I also need the axis of the graph of function for:
f(x)=4x^2-32x+27
complete the square
y=4(x^2-8x+16)+27-64
y=4(x-4)^2-37
vertex: (4,-37)
axis of symmetry: x=4 (parabola opens upwards)
..
As well as the quadratic function in vertex form of:
y=x^2+8x
complete the square
y=(x^2+8x+16)-16
y=(x+4)^2-16
vertex: (-4,-16)